Solve for the following equation: 7.701-y=62+149.81-173 I would greatly appreciate your assistance in this equation.
Answer by rothauserc(4718)

Combine like terms
7.701 - y = 38.81
-y = 31.109
Divide both sides of = by -
y = -31.109